# What Does ‘Human-Centred AI’ Mean?

### v1

@@ 1  Introduction
The process of uncondensing time indubitably requires slowing down [281] and radically centring the human and decentring the “ultra rapid computing machine.” [201] To do this, some propose to (conceptually) reengineer [244] or reimagine these machines and our relationships to them inter alia through a subfield known as human-centred AI. For a pertinent example, Andy J. [305], in the call for this special issue, describes human-centred artificial intelligence (HCAI) as placing “human behaviour and experience at the heart of [AI] research.” An example of this, he says, could be cases wherein it is “claimed that artificial neural networks (ANNs) now perform at human levels in a variety of tasks”. He goes on to ask:
@@
Overall we see that HCAI — as a specialisation for students [214, 216, e.g.], as a subfield for scholars [305, 212, 267, 215, 203, 279, e.g.] or technology industry workers [188, e.g.] alike, and as a concern and mission for regulators [256, 286, e.g.] — interweaves four related themes:
@@
In this paper, I turn this frame on its head [161, 267, 268, 272, 247, 297, 159, 277, 170, 280, cf.]. Using a unique (re)definition of AI that releases us from a correlationist [^1] The idea that it is correlations with a given target, like human data derived benchmarks ([273, 251]; cf. [296]), that validate the safety [184], appropriateness, or other desirable properties of a model or system. In general, correlationism may be unproblematic, but in a setting where computation is involved correlations cannot function as a useful guide, serving more as red herring than anything else [196, 197, 195, 194]. grip, we then examine three case studies of techno-social relationships between cognition and artifacts. To presage the coming analyses, herein AI is any techno-social relationship that outsources to machines or algorithms some part, however small, of human cognitive labour [296, cf.]. I demonstrate that AI is human-centric, not because it behaves like or is designed to be like humans, but because it requires a ghost in the machine, often literally an obfuscated human-in-the-loop to properly function [196, also see] because AI is humans albeit in fetishised [241, 172, 243], obfuscated forms [185, e.g.]. That is, AI “is in reality produced by relations among people [even though it] appears before us in a fantastic form as relations among things” [254]. AI’s “technological veil” hides human cognitive labour [243]. I bring this anthropological, sociotechnical and broadly computational cognitive scientific angle to understanding AI, that is unlike “the skills and techniques of behavioural research, previously applied to humans and other animals[, which claim to help us] better understand the ‘psychology’ of complex black-box ANNs” ([305]; cf. [262]). In fact behavioural probing of such systems using such experimental techniques assumes their psychological standing to be equal to, or comparable to, biological organisms: it begs the question [195, 194, 186, 261, see for relevant analyses:].
@@ 2  Radically Redefining AI
As practitioners of science, we are duty bound to think about what ‘artificial’ and ‘intelligence’ mean. What does it mean to propose an artificial version of a human capacity? When we talk about redefining something, we must remember that “definitions have no inherent truth. They are agreed conventions tested by their internal coherence as well as by their relationship to common sense, common practice and history.” ([310]; also see [183, 244]) Some of the existing definitions of AI when probed reveal weaknesses in the aforementioned dimensions — especially undesirable when speaking of formal systems, which all AI systems necessarily are. This is to say, “weakness from the viewpoint of formal symmetry of doctrine [is] strength in the service of rising capitalism.” [248, p. 25] So these formal and common-sense problems with pre-existing AI conceptions form a core of strength when incentives are, in opposition to academia, ones of profit-making for the technology industry (see [296], for an exposition of possible meanings). AI has had so many related but nonetheless different meanings over the decades that perhaps it is a fool’s errand to even try and pin it down [195, 171, 225, 200, 179, 234, 278]. And similar issues with terminology appear with related terms like ‘brain inspired’ and ‘neurally plausible’ to name but a few terms which function as weasel words [195] — as well as when ‘computational’ is used metaphorically or naively by cognitive neuroscientists [196, 197].
@@
Without further ado and notwithstanding terminological disarray, Table 1 embodies a single tripartite definition that acts as an analytical tool for discerning the AI’s properties, e.g. labour obfuscation, which is the extent of the hidden human-in-the-loop [196, 176]. The radical redefinition I propose comprises two parts. First, full-blown deflation — any sociotechnical relationship could be AI if it links an artefact to human cognition such that the artefact can be seen as performing some aspect of a cognitive capacity. For this step, all we need to ask ourselves is: does the technology exist in a relationship with human cognition? If we affirm this, we can move to the next step. Second, we reinflate AI into three types of relationship, that of enhancement, of replacement, and of displacement (see Table 1) of human cognitive labour.
@@ 3  Artefacts versus Cognitive Labour
AI is unlike a tool like a saw used to cut wood where the person cutting also puts in labour, often more than the creator of the saw, to cut wood, in the sense that a saw is involved in the predominantly overt physical and cognitive labour of woodworking [198]. Not so for typical cases of contemporary AI, like chatbots — in contrast: You did not contribute other than as input, e.g. the so-called prompt, ultimately harming the chance to learn anything substantial [163, 198, e.g.]. It is never clear if the chatbot’s results really match those desired, e.g. so-called hallucinations — “a misleading (and anthropomorphizing) term[, which has become mainstream, and thus] a major win for AI hype” ([207], p. 700; [209, 170]). Due to their opaque nature, AI relationships hide labour that might be invoked not only like the saw to extract raw materials or design systems, but also in real-time as you use the artefact, e.g. you forget that sweatshop workers may be in real-time or the recent past refining the output of the chatbot [176, 252, 257]. So much for a stark contrast between a tool-human sociotechnical relationship without AI and a relationship that involves modern chatbots, what about technologies that are more complex than a saw?
@@
: these three (in Table 4) implicate what is typical contemporary AI, very deep artificial neural networks trained on extremely large datasets [195, a correlationist programme previously dubbed modern connectionism in] with the cognitive labour they claim to capture, of which all relationships are characterised as displacement.
@@ 3.1 Abacus, Calculator, Computer
The same holds for the human computers in the USA at the National Aeronautics and Space Administration [276, NASA;]:
@@
But when they went to computers, they called over and said, “tell her to check and see if the computer trajectory they had calculated was correct.” So I checked it and it was correct. [211, Katherine Johnson, as quoted in]
@@ 3.2 Alarm Clock, Camera, Garment Factory
The camera is perhaps a more understandable addition to Table 3 with its relationship to the human capacity of vision well-known. The point here — as with all the other relationships — is not a mechanistic similarity but one based on functional correspondence or role, factual or perceived [195, 194, 197, cf.]. Notable in our relationship with the camera is the fact it has largely deskilled nothing in the present for the typical user, using a camera or looking at its output, does not negatively impact our ability to visually perceive (rows b & c, Table 3). Photographs also serve, especially in the advent of mobile phone cameras, to help us recollect our memories, enhancing our own ability to think about the past (row a). Of course, the user, the photographer, must frame and perform automatic or manual adjustments to the lens, for any image or video to be of use (row g). However, this human contribution, which manifests as the user, is ever present in all our AI systems from calculators to alarm clocks to much more complex systems like digital computers. The more infrequently discussed human-in-the-loop is humans other than the user, such as cases like digital computers (3rd column of Table 2) which require many other people’s labour, not just the user’s [221, 169, 220, 166, 168, 175, 176, 257]. An infamous example of the human-in-the-loop technique is the orientalist Mechanical Turk, which toured from the late 18th to the mid 19th century, wherein a person hid in a cabinet under what appeared to be an automaton that played chess [282]. In fact, the person below the puppet controlled its movements, giving the human player sitting across it the impression that they were being beaten at chess by a clockwork machine. This is also the namesake of Amazon’s Mechanical Turk, a platform on which low paid workers toil to “earn pennies or dollars doing tasks that computers cannot yet [and may never] easily do.” [245]
@@
The sociotechnical relationship between the garment factory, which comprises humans-in-the-loop, can be seen as a more harmful, and equally obfuscatory, variant of the Mechanical Turk. Workers are treated badly in many cases, such as where sweatshop labour is used, the environmental impact of so many clothes is ignored, and the harmful chemicals are glossed over. In the not so distant past, people owned fewer clothes and had them hand- and custom-made by seamstress and tailors. We readily have accepted this displacement of a relationship that involved a skilled adult making us a small selection of well-fitting clothes, which typically could last a whole lifetime, to a world in which we consume clothes relentlessly made often by workers, who can even be underage, and are often in harmful conditions [247, 252, 282, 176, 301, 280, cf.]. Notwithstanding, bespoke tailoring is to this day understood to be superior, because it factually is, and so preferred by the rich and famous and indeed required by anybody outside the bounds of factory-made standardised clothing configurations and sizes (row e). Even the simpler skill (compared to sewing from scratch) of taking in or out clothes as our bodies change over time is abandoned (row c) as a function of the garment factory which produces new clothes cheaply that we can buy instead (row e). We now turn to the cutting edge of contemporary AI.
@@ 3.3 LLM, Image Generator, Chatbot
If we centre the cognitive labour when teasing apart these three AI relations, as Table 1 guides us to do, it falls into our lap how it is displacement, harmful, and deskilling in every case: large language model (LLMs) versus essay writing, image generator versus artist, and chatbot versus human companionship (rows ac, Table 4). Importantly, as [263] explain, “whereas in human-animal interaction there is ample evidence of reciprocal adaptation, here the adaptation is strikingly one-sided, with [such systems] essentially helpless, requiring care [227] and forcing people to adapt to their constraints [158, 285].” We are expected somehow to learn what is presented as a non-skill, because ‘prompt engineering’ is indeed not a skill, to enable us to use these opaque corporate-owned stochastic context-addressable systems. And, as turbo charged versions of the Mechanical Turk, contemporary AI models “are built on massive exploitative ‘ghost labour’; crowdsourced and outsourced labour that follows the patterns of colonial relations [164, e.g.,][236].
The false advertising — there is no engineering in prompt engineering — grows the longer the rhetoric around these systems are examined because, as we have seen many times so far, “data can only do so much. [In any apparently successful application of AI t]he real work is carried out by the people on the ground” [178]. While in the previous century, “computing systems functioned due to vast arrays of human workers, expressed through workflow organization, operators’ actions, and software” [208], in the present the obfuscated human-in-the-loop — such as the sweatshop workers who guide and power LLMs and other such systems, or us ourselves whose data is stolen without our knowledge — is not respected [250, 185, 235, 167, 192]. Even the low bar human computers’ treatment has set is not met by modern AI’s dehumanisation and theft of labour [294, 173, 176, 252, 189, 265, e.g.]. In the time of software updates, the obfuscation of programmers’ labour with seamless updates grows, while in the past one would physically go get artifacts fixed (or indeed fix them ourselves).
@@ 4  Machine Hauntology & Spectral Technology
    We can centre human cognition, and therefore we can centre the study thereof, cognitive science, as a relevant discipline to understand purported cases of artificial cognition [296, e.g.]. Under computationalism, which the possibility of engineering cognition implicates, correlations and benchmarks are not relevant [197, 196, e.g.]. This allows us to in principle reject any argument that uses behavioural or neuroimaging correlation to argue for human-likeness of an artifact.
@@
    We can reject AI hype, anthropomorphism, mysterianism about known mechanism, [^4] The idea that engineered mechanisms, like matrix multiplication or other operations in artificial neural networks are somehow uniquely unknown or unknowable properties of these systems. exaggeration, or fads, removing this rhetoric from being relevant as to what counts as AI, which is a typical frame with many definitions of AI, often to push products [186, 180, e.g.]. The same goes for claims about neural or biological plausibility or inspiration; these have no useful coherent definitions [195, 194, e.g.].
@@
    We can consider AI in general abstracted terms without requiring specific reference to current advances in AI, e.g. chess-playing algorithms are AI, regardless of whether systems are cutting edge or not [234, sometimes called the AI effect;]. And so we can easily reject that only artificial neural nets or large language models or generative AI are AI, especially when until recently GOFAI (good old-fashioned AI, also known as symbolic) was canonically AI, hence the name [198].
@@
    Relatedly, we can grant AI a (pre)history, allowing us to include the Antikythera mechanism [187], astrolabes, sextants, abacuses, and more in the timeline of AI [185, 232, e.g.]. We can uncondense time — allowing us to slow down and giving us back our history — which is centrally relevant for understanding our present or possible futures [201, 281]. As mentioned, this is something of a phobia, notably:
@@
And the same for other so-called founding fathers of AI, who when coining ‘Artificial Intelligence’ claimed “that every aspect of learning or any other feature of intelligence can in principle be so precisely described that a machine can be made to simulate it” [233]. This foundational document of AI, which was a proposal for a two month long summer conference, shows all the hallmarks we wish to avoid. As described by [253]: “The starting point for this multidisciplinary effort was the coupling of ‘natural and artificial intelligence,’ although, as the Dartmouth proposal made clear, the distinction between the two was uninteresting for many. These two realms constituted a common problem.” In a now classic act of condensing time [281, 201, recall], [233] infamously proposed that this project of ‘solving’ intelligence would take a single season. Not in this AI summer, and not in any, will mathematically impossible displacements of humans pass muster [296, 266].
@@
I do not say the above to motivate rejections of mainstream computationalism, but if the naive conception furthers mystifying basic computational processes and engineering then it surely must, we must make it, give way to more mature ways of thought that take into account what it is we commit to [196, cf. basic versus naive versus non-naive computationalisms,]. In other words, just because the destination along the road is currently invisible to the naked eye from the origin does not by any stretch of the imagination imply something magical will get us there. What gets us to the destination — any goal — is our pre-existing cognition whether the end is currently visible to the naked eye or not. When it comes to the part of getting there that is automated, from car to calculator, the systems’ parts are known and just hum along similarly regardless of whether the distance is 50 meters or 50 thousand or whether the digits are in the tens or the trillions.

### v2

@@ 1  Introduction
The process of uncondensing time indubitably requires slowing down [285] and radically centring the human and decentring the “ultra rapid computing machine.” [205] To do this, some propose to (conceptually) reengineer [248, 195] or reimagine these machines and our relationships to them inter alia through a subfield known as human-centred AI. For a pertinent example, Andy J. [309], in the call for this special issue, describes human-centred artificial intelligence (HCAI) as placing “human behaviour and experience at the heart of [AI] research.” An example of this, he says, could be cases wherein it is “claimed that artificial neural networks (ANNs) now perform at human levels in a variety of tasks”. He goes on to ask:
@@
Overall we see that HCAI — as a specialisation for students [218, 220, e.g.], as a subfield for scholars [309, 216, 271, 219, 207, 283, e.g.] or technology industry workers [191, e.g.] alike, and as a concern and mission for regulators [260, 290, e.g.] — interweaves four related themes:
@@
In this paper, I turn this frame on its head [163, 271, 272, 276, 251, 301, 161, 281, 172, 284, cf.]. Using a unique (re)definition of AI that releases us from a correlationist [^1] The idea that it is correlations with a given target, like human data derived benchmarks ([277, 255]; cf. [300]), that validate the safety [187], appropriateness, or other desirable properties of a model or system. In general, correlationism may be unproblematic, but in a setting where computation is involved correlations cannot function as a useful guide, serving more as red herring than anything else [200, 201, 199, 198]. grip, we then examine three triplets as case studies of techno-social relationships between cognition and artifacts. To presage the coming analyses, herein AI is any techno-social relationship that outsources to machines or algorithms some part, however small, of human cognitive labour [300, cf.]. I demonstrate that AI is human-centric, not because it behaves like or is designed to be like humans, but because it requires a ghost in the machine, often literally an obfuscated human-in-the-loop to properly function [200, also see] because AI is humans albeit in fetishised [245, 174, 247], obfuscated forms [188, e.g.]. That is, AI “is in reality produced by relations among people [even though it] appears before us in a fantastic form as relations among things” [258]. AI’s “technological veil” hides human cognitive labour [247]. I bring this anthropological, sociotechnical and broadly computational cognitive scientific angle to understanding AI, that is unlike “the skills and techniques of behavioural research, previously applied to humans and other animals[, which claim to help us] better understand the ‘psychology’ of complex black-box ANNs” ([309]; cf. [266]). In fact behavioural probing of such systems using such experimental techniques assumes their psychological standing to be equal to, or comparable to, biological organisms: it begs the question [199, 198, 189, 265, see for relevant analyses:].
@@ 2  Radically Redefining AI
As practitioners of science, we are duty bound to think about what ‘artificial’ and ‘intelligence’ mean. What does it mean to propose an artificial version of a human capacity? When we talk about redefining something, we must remember that “definitions have no inherent truth. They are agreed conventions tested by their internal coherence as well as by their relationship to common sense, common practice and history.” ([314]; also see [186, 248]) Some of the existing definitions of AI when probed reveal weaknesses in the aforementioned dimensions — especially undesirable when speaking of formal systems, which all AI systems necessarily are. This is to say, “weakness from the viewpoint of formal symmetry of doctrine [is] strength in the service of rising capitalism.” [252, p. 25] So these formal and common-sense problems with pre-existing AI conceptions form a core of strength when incentives are, in opposition to academia, ones of profit-making for the technology industry (see [300], for an exposition of possible meanings). AI has had so many related but nonetheless different meanings over the decades that perhaps it is a fool’s errand to even try and pin it down [199, 173, 229, 204, 181, 238, 282]. And similar issues with terminology appear with related terms like ‘brain inspired’ and ‘neurally plausible’ to name but a few terms which function as weasel words [199] — as well as when ‘computational’ is used metaphorically or naively by cognitive neuroscientists [200, 201].
@@
Without further ado and notwithstanding terminological disarray, Table 1 embodies a single tripartite definition that acts as an analytical tool for discerning the AI’s properties, e.g. labour obfuscation, which is the extent of the hidden human-in-the-loop [200, 178]. The radical redefinition I propose comprises two parts. First, full-blown deflation — any sociotechnical relationship could be AI if it links an artefact to human cognition such that the artefact can be seen as performing some aspect of a cognitive capacity [[, cf.]]egan2025deflating, pittphilsci24834. For this step, all we need to ask ourselves is: does the technology exist in a relationship with human cognition? If we affirm this, we can move to the next step. Second, we reinflate AI into three types of relationship, that of enhancement, of replacement, and of displacement (see Table 1) of human cognitive labour.
@@ 3  Artefacts versus Cognitive Labour
AI is unlike a tool like a saw used to cut wood where the person cutting also puts in labour, often more than the creator of the saw, to cut wood, in the sense that a saw is involved in the predominantly overt physical and cognitive labour of woodworking [202]. Not so for typical cases of contemporary AI, like chatbots — in contrast: You did not contribute other than as input, e.g. the so-called prompt, ultimately harming the chance to learn anything substantial [165, 202, e.g.]. It is never clear if the chatbot’s results really match those desired, e.g. so-called hallucinations — “a misleading (and anthropomorphizing) term[, which has become mainstream, and thus] a major win for AI hype” ([211], p. 700; [213, 172]). Due to their opaque nature, AI relationships hide labour that might be invoked not only like the saw to extract raw materials or design systems, but also in real-time as you use the artefact, e.g. you forget that sweatshop workers may be in real-time or the recent past refining the output of the chatbot [178, 256, 261]. So much for a stark contrast between a tool-human sociotechnical relationship without AI and a relationship that involves modern chatbots, what about technologies that are more complex than a saw?
@@
: these three (in Table 4) implicate what is typical contemporary AI, very deep artificial neural networks trained on extremely large datasets [199, a correlationist programme previously dubbed modern connectionism in] with the cognitive labour they claim to capture, of which all relationships are characterised as displacement.
@@ 3.1 Abacus, Calculator, Computer
The same holds for the human computers in the USA at the National Aeronautics and Space Administration [280, NASA;]:
@@
But when they went to computers, they called over and said, “tell her to check and see if the computer trajectory they had calculated was correct.” So I checked it and it was correct. [215, Katherine Johnson, as quoted in]
@@ 3.2 Alarm Clock, Camera, Garment Factory
The camera is perhaps a more understandable addition to Table 3 with its relationship to the human capacity of vision well-known. The point here — as with all the other relationships — is not a mechanistic similarity but one based on functional correspondence or role, factual or perceived [199, 198, 201, cf.]. Notable in our relationship with the camera is the fact it has largely deskilled nothing in the present for the typical user, using a camera or looking at its output, does not negatively impact our ability to visually perceive (rows b & c, Table 3). Photographs also serve, especially in the advent of mobile phone cameras, to help us recollect our memories, enhancing our own ability to think about the past (row a). Of course, the user, the photographer, must frame and perform automatic or manual adjustments to the lens, for any image or video to be of use (row g). However, this human contribution, which manifests as the user, is ever present in all our AI systems from calculators to alarm clocks to much more complex systems like digital computers. The more infrequently discussed human-in-the-loop is humans other than the user, such as cases like digital computers (3rd column of Table 2) which require many other people’s labour, not just the user’s [225, 171, 224, 168, 170, 177, 178, 261]. An infamous example of the human-in-the-loop technique is the orientalist Mechanical Turk, which toured from the late 18th to the mid 19th century, wherein a person hid in a cabinet under what appeared to be an automaton that played chess [286]. In fact, the person below the puppet controlled its movements, giving the human player sitting across it the impression that they were being beaten at chess by a clockwork machine. This is also the namesake of Amazon’s Mechanical Turk, a platform on which low paid workers toil to “earn pennies or dollars doing tasks that computers cannot yet [and may never] easily do.” [249]
@@
The sociotechnical relationship between the garment factory, which comprises humans-in-the-loop, can be seen as a more harmful, and equally obfuscatory, variant of the Mechanical Turk. Workers are treated badly in many cases, such as where sweatshop labour is used, the environmental impact of so many clothes is ignored, and the harmful chemicals are glossed over. In the not so distant past, people owned fewer clothes and had them hand- and custom-made by seamstresses and tailors. We readily have accepted this displacement of a relationship that involved a skilled adult making us a small selection of well-fitting clothes, which typically could last a whole lifetime, to a world in which we consume clothes relentlessly made often by workers, who can even be underage, and are often in harmful conditions [251, 256, 286, 178, 305, 284, cf.]. Notwithstanding, bespoke tailoring is to this day understood to be superior, because it factually is, and so preferred by the rich and famous and indeed required by anybody outside the bounds of factory-made standardised clothing configurations and sizes (row e). Even the simpler skill (compared to sewing from scratch) of taking in or out clothes as our bodies change over time is abandoned (row c) as a function of the garment factory which produces new clothes cheaply that we can buy instead (row e). We now turn to the cutting edge of contemporary AI.
@@ 3.3 LLM, Image Generator, Chatbot
If we centre the cognitive labour when teasing apart these three AI relations, as Table 1 guides us to do, it falls into our lap how it is displacement, harmful, and deskilling in every case: large language model (LLMs) versus essay writing, image generator versus artist, and chatbot versus human companionship (rows ac, Table 4). Importantly, as [267] explain, “whereas in human-animal interaction there is ample evidence of reciprocal adaptation, here the adaptation is strikingly one-sided, with [such systems] essentially helpless, requiring care [231] and forcing people to adapt to their constraints [160, 289].” We are expected somehow to learn what is presented as a non-skill, because ‘prompt engineering’ is indeed not a skill, to enable us to use these opaque corporate-owned stochastic context-addressable systems. And, as turbo charged versions of the Mechanical Turk, contemporary AI models “are built on massive exploitative ‘ghost labour’; crowdsourced and outsourced labour that follows the patterns of colonial relations [166, e.g.,][240].
The false advertising — there is no engineering in prompt engineering — grows the longer the rhetoric around these systems are examined because, as we have seen many times so far, “data can only do so much. [In any apparently successful application of AI t]he real work is carried out by the people on the ground” [180]. While in the previous century, “computing systems functioned due to vast arrays of human workers, expressed through workflow organization, operators’ actions, and software” [212], in the present the obfuscated human-in-the-loop — such as the sweatshop workers who guide and power LLMs and other such systems, or us ourselves whose data is stolen without our knowledge — is not respected [254, 188, 239, 169, 196]. Even the low bar human computers’ treatment has set is not met by modern AI’s dehumanisation and theft of labour [298, 175, 178, 256, 192, 269, e.g.]. In the time of software updates, the obfuscation of programmers’ labour with seamless updates grows, while in the past one would physically go get artifacts fixed (or indeed fix them ourselves).
@@ 4  Machine Hauntology & Spectral Technology
    We can centre human cognition, and therefore we can centre the study thereof, cognitive science, as a relevant discipline to understand purported cases of artificial cognition [300, e.g.]. Under computationalism, which the possibility of engineering cognition implicates, correlations and benchmarks are not relevant [201, 200, e.g.]. This allows us to in principle reject any argument that uses behavioural or neuroimaging correlation to argue for human-likeness of an artifact.
@@
    We can reject AI hype, anthropomorphism, mysterianism about known mechanism, [^4] The idea that engineered mechanisms, like matrix multiplication or other operations in artificial neural networks are somehow uniquely unknown or unknowable properties of these systems. exaggeration, or fads, removing this rhetoric from being relevant as to what counts as AI, which is a typical frame with many definitions of AI, often to push products [189, 182, e.g.]. The same goes for claims about neural or biological plausibility or inspiration; these have no useful coherent definitions [199, 198, e.g.].
@@
    We can consider AI in general abstracted terms without requiring specific reference to current advances in AI, e.g. chess-playing algorithms are AI, regardless of whether systems are cutting edge or not [238, sometimes called the AI effect;]. And so we can easily reject that only artificial neural nets or large language models or generative AI are AI, especially when until recently GOFAI (good old-fashioned AI, also known as symbolic) was canonically AI, hence the name [202].
@@
    Relatedly, we can grant AI a (pre)history, allowing us to include the Antikythera mechanism [190], astrolabes, sextants, abacuses, and more in the timeline of AI [188, 236, e.g.]. We can uncondense time — allowing us to slow down and giving us back our history — which is centrally relevant for understanding our present or possible futures [205, 285]. As mentioned, this is something of a phobia, notably:
@@
And the same for other so-called founding fathers of AI, who when coining ‘Artificial Intelligence’ claimed “that every aspect of learning or any other feature of intelligence can in principle be so precisely described that a machine can be made to simulate it” [237]. This foundational document of AI, which was a proposal for a two month long summer conference, shows all the hallmarks we wish to avoid. As described by [257]: “The starting point for this multidisciplinary effort was the coupling of ‘natural and artificial intelligence,’ although, as the Dartmouth proposal made clear, the distinction between the two was uninteresting for many. These two realms constituted a common problem.” In a now classic act of condensing time [285, 205, recall], [237] infamously proposed that this project of ‘solving’ intelligence would take a single season. Not in this AI summer, and not in any, will mathematically impossible displacements of humans pass muster [300, 270].
@@
I do not say the above to motivate rejections of mainstream computationalism, but if the naive conception furthers mystifying basic computational processes and engineering then it surely must, we must make it, give way to more mature ways of thought that take into account what it is we commit to [200, cf. basic versus naive versus non-naive computationalisms,]. In other words, just because the destination along the road is currently invisible to the naked eye from the origin does not by any stretch of the imagination imply something magical will get us there. What gets us to the destination — any goal — is our pre-existing cognition whether the end is currently visible to the naked eye or not. When it comes to the part of getting there that is automated, from car to calculator, the systems’ parts are known and just hum along similarly regardless of whether the distance is 50 meters or 50 thousand or whether the digits are in the tens or the trillions.